# Unveiling the Sagittarius Dwarf Spheroidal Galaxy Core with Gaia DR3

### v1

@@ 4 Red Clump Distance Determination
We use a Monte Carlo method to determine the uncertainties on our distance moduli and their corresponding distances. For each distribution of for the Sgr core and M54, we generated a KDE with a bandwidth value of 0.04. From our KDEs, we generated 10,000 resampled distributions over the range our RC was located, ending with a distribution of apparent magnitudes () for our objects. We then modeled the extinction values , generating a distribution based on our results from Bayestar19, resampling 10,000 times. From the absolute magnitude of the RC in the 2MASS band from Hawkins et al. (2017), we generate a normal distribution with a mean of mag and a standard deviation of mag. We calculate the distance moduli for the Sgr core and M54 using these distributions, along with their heliocentric distances, extracting the 16th, 50th, and 84th percentiles to provide the distance moduli to the Sgr core and M54 with errors, along with their distances.
@@ 5 The Connection of the Sgr Core and M54
We isolated the Sgr core and Messier 54 spatially using DR3 astrometry, applied general quality cuts from Gaia Collaboration et al. (2021), and cleaned Milky Way contaminants from our sample by identifying member stars within the histograms of proper motion in right ascension direction and declination direction . We also calibrated distances to both of these systems using a red clump distance determination. Both distance moduli found using red clump distance methods of mag and mag fall well within literature values from Mateo et al. (1995); Monaco et al. (2004); Kunder & Chaboyer (2009); Sollima et al. (2010), as shown in Figure 4, and prove both the precision of our sample and effectiveness of red clump distance methods. Our measured distance moduli to both systems and their respective errors overlap, suggesting that the Sgr core and M54 could occupy the same space. However, as this method only assigns one distance to the entire system–more analysis is needed. A future look into individual star distance determination methods would provide conclusive evidence if M54 is projected onto the area of Sgr’s nucleus.
@@
We corroborate our membership selection using available APOGEE DR17 metallicities and radial velocities from our member stars from the Sgr core and M54. From our Sgr core member selection, we have 1489 stars that have radial velocities and metallicities, and from M54 we have a total of 97 member stars. We show the distribution of radial velocities and metallicities in Figure 5.
We report a range of [Fe/H] of to dex for the Sgr core, with a mean radial velocity of . For M54, we report a [Fe/H] range of to dex, with a mean radial velocity of . Visually, the radial velocities for the Sgr core and M54 are similar, with only difference in their mean. Meanwhile, the Sgr core has a wide-range of [Fe/H], with a high of nearly solar [Fe/H], to a low of dex, matching in the range found in Siegel et al. (2007); Alfaro-Cuello et al. (2019). M54 has a binomial distribution of [Fe/H], matching the findings in Minelli et al. (2023). We find a smaller range of [Fe/H] within M54 with a dex range difference from the Sgr core. We were unable to perform a comparison to the found one dex difference of median [Fe/H] as found in Hayes et al. (2020) did in the streams, due to the slight bimodal distribution featured in both the Sgr core and M54.
@@
The stellar density contours present in the bottom panel of Figure 1 instead showcase a concentration in the central area of the Sgr core near M54, implying decoupling between the motion of the Sgr core its central concentration peak, possibly indicating that M54 remains untouched despite being embedded in the Sgr core.
@@
Evidence supporting the theory that M54 formed independently of the Sgr core comes from their respective red giant branch from their CMDs made from our sample in Figure 3. In the bottom panels, where the RGB and AGB branches are indicated with APOGEE DR17 [Fe/H], the Sgr core displays a rough gradient along its RGB and corresponding AGB branches, marked by three different metallicities. This lines up with the three pericentric interactions with the core of the Milky Way that kick-started star formation in both galaxies (de Boer et al., 2015; Ruiz-Lara et al., 2020). The CMD of M54 lacks these three distinct RGB and AGB branches, instead bearing two distinct branches as found in Figure 2 of Mucciarelli et al. (2017). This suggests that M54 only went through one or two pericentric passages, enriching the cluster, kick-starting star formation, implying M54 infell during the first passage of Sgr. Similarly, the density contours present in the bottom right panel of Figure 1 and the jump from the uniform center to a scattered value outside the first density contour lines with the central density cusp found in Monaco et al. (2005) validate our sample further.
@@
The most recent paper for a Sgr core membership selection was Minelli et al. (2023), where they identified 450 member stars in Early Data Release 3 (EDR3). Cross-matching with our sample, we recover 427 stars from Minelli’s sample. From those 427 stars, none have [Fe/H] measurements from APOGEE DR17, and there are no DR3 [M/H] measurements using General Stellar Parametrizer from Spectroscopy (GSP-Spec) for our sample. There are [M/H] measurements for these 427 within DR3 General Stellar Parametrizer from Photometry (GSP-Phot), however these values exhibit substantial biases compared to literature values and are only useful at a qualitative level especially for red giant branch stars with low quality parallax measurements (Andrae et al., 2023). Other attempts to gain [Fe/H] measurements from The Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) and The GALactic Archaeology with HERMES (GALAH) fail to capture Sgr dSph core and M54 member stars within our sample, unfortunately not allowing us a direct comparison of Minelli et al., 2023 [Fe/H] measurements taken from the Fibre Large Array Multi Element Spectrograph (FLAMES) on the Very Large Telescope (VLT).
6 Conclusion
@@
We demonstrate the strength of DR3 astrometry combined with APOGEE DR17 spectra to separate the Sgr core with M54, along with a multi-Gaussian fit to reduce MW contaminants for systems near the disk. We have a membership catalog of 515689 stars in the Sgr core and 4512 in M54.
We measure the distance to both the Sgr core and M54 using red clump distancing methods from a cross-matched sample with 2MASS, finding a heliocentric distance of kpc for the Sgr Core and kpc for M54. Our distance measurements agree with the literature but are more precise due to the larger member sample identified in our work.
We analyze the motion, density contours, and [Fe/H] of the Sgr core and M54 to understand whether M54 formed independently and was captured during a passage, or is the Sgr nucleus. Preliminary distance measurements and the dynamics of our sample imply they co-exist, but the chemical abundances suggest the systems formed independently.
@@ 7 Acknowledgments
E.K.H.T.T. would like to additionally extend gratitude to Dr. Robyn Sanderson from the University of Pennsylvania, Dr. Sukanya Chakrabarti from the University of Alabama, Huntsville, Dr. Marc Hon from the Massachusetts Institute of Technology, Dr. Andrea Dobson from Whitman College, Dr. Matthew Shetrone from the University of Californa, Santa Cruz, and Aldo Sepuvelda from the University of Texas at Austin School of Law for their insightful conversations and feedback on this work.
@@
8 Appendix
8.1 Gaia DR3 ADQL Queries
The following DR AQDL queries produce our spatial sample for the Sgr dSph core and Milky Way contamination, respectively, in full.
SELECT TOP 50000000
@@ 8.1 Gaia DR3 ADQL Queries
SELECT TOP 50000000 gaia_source.designation, gaia_source.source_id, gaia_source.ra, gaia_source.ra_error, gaia_source.dec, gaia_source.dec_error, gaia_source.parallax, gaia_source.parallax_error, gaia_source.parallax_over_error, gaia_source.pm, gaia_source.pmra, gaia_source.pmra_error, gaia_source.pmdec, gaia_source.pmdec_error, gaia_source.visibility_periods_used, gaia_source.phot_g_mean_mag, gaia_source.bp_rp, gaia_source.radial_velocity, gaia_source.radial_velocity_error, gaia_source.rv_template_fe_h, gaia_source.phot_variable_flag, gaia_source.l, gaia_source.b, gaia_source.non_single_star, gaia_source.has_epoch_rv, gaia_source.teff_gspphot, gaia_source.logg_gspphot, gaia_source.mh_gspphot, gaia_source.distance_gspphot, gaia_source.ag_gspphot, gaia_source.ebpminrp_gspphot FROM gaiadr3.gaia_source WHERE CONTAINS( POINT(’ICRS’, gaiadr3.gaia_source.ra, gaiadr3.gaia_source.dec), CIRCLE(’ICRS’, 259.65, -18.47, 3))=1 AND gaia_source.phot_g_mean_mag IS NOT NULL AND gaia_source.parallax IS NOT NULL AND gaia_source.bp_rp IS NOT NULL
8.2 Flattened Coordinates
Below we showcase the equations to compute the flattened coordinates for the Sgr core and account of the elliptical shape of the core.
8.3 Candidate Members Table
Table 4 and Table 5 (full tables available at CDS), contain the data for identified members of the Sgr core and Messier 54. Below we show an excerpt of the first five rows from each system.
| Gaia Source_id | ||||||
| 6736253232769633408 | 282.81 | -31.56 | -2.41 | -1.36 | 17.96 | 1.24 |
| 6736253542006710528 | 282.80 | -31.54 | -2.91 | -1.07 | 17.92 | 1.37 |
| 6736253885604674688 | 282.72 | -31.54 | -2.50 | -1.43 | 17.82 | 1.29 |
| 6736254126121627648 | 282.68 | -31.53 | -2.57 | -1.12 | 17.97 | 1.29 |
| 6736254233517136128 | 282.75 | -31.54 | -0.57 | -2.72 | 17.91 | 1.12 |
Full table is available at the CDS.
| Gaia DR3 Source_id | ||||||
| 6760412629965372928 | 283.83 | -30.56 | -2.89 | -1.52 | 17.97 | 1.40 |
| 6760412629966580096 | 283.84 | -30.56 | -2.70 | -1.08 | 17.84 | 1.26 |
| 6760412629969429632 | 283.83 | -30.57 | -2.79 | -1.07 | 17.94 | 1.29 |
| 6760412634271816448 | 283.84 | -30.57 | -2.65 | -1.70 | 17.96 | 1.30 |
| 6760412634289471488 | 283.84 | -30.57 | -2.56 | -1.40 | 18.01 | 1.40 |
Full table is available at the CDS.

### v2

@@ IV Red Clump Distance Determination
We employ a Monte Carlo method to determine the uncertainties on our distance moduli and their corresponding distances. For each distribution of for the Sgr core and M54, we generated a KDE with a bandwidth value of 0.04. From our KDEs, we generated 10,000 resampled distributions over the range our RC was located, ending with a distribution of apparent magnitudes () for our objects. We then modeled the extinction values , generating a distribution based on our results from Bayestar19, resampling 10,000 times. From the absolute magnitude of the RC in the 2MASS band from Hawkins et al. (2017), we generate a normal distribution with a mean of mag and a standard deviation of mag. We calculate the distance moduli for the Sgr core and M54 using these distributions, along with their heliocentric distances, extracting the 16th, 50th, and 84th percentiles to provide the distance moduli to the Sgr core and M54 with errors, along with their distances.
@@ V The Connection of the Sgr Core and M54
We isolated the Sgr core and Messier 54 spatially using DR3 astrometry, applied general quality cuts from Gaia Collaboration et al. (2021), and cleaned Milky Way contaminants from our sample by identifying member stars within the histograms of proper motion in right ascension direction and declination direction . We also calibrated distances to both of these systems using a red clump distance determination. Both distance moduli found using red clump distance methods of mag and mag fall well within literature values from Mateo et al. (1995); Alard (1996); Monaco et al. (2004); Siegel et al. (2007); Kunder & Chaboyer (2009); Sollima et al. (2010), as shown in Figure 4, and prove both the precision of our sample and effectiveness of red clump distance methods. Our measured distance moduli to both systems and their respective errors overlap, suggesting that the Sgr core and M54 could occupy the same space. However, as this method only assigns one distance to the entire system–more analysis is needed. A future look into individual star distance determination methods would provide conclusive evidence if M54 is projected onto the area of Sgr’s nucleus.
@@
We corroborate our membership selection using available APOGEE DR17 metallicities and radial velocities from our member stars from the Sgr core and M54. From our Sgr core member selection, we have 1589 stars that have radial velocities and metallicities, and from M54 we have a total of 104 member stars. We show the distribution of radial velocities and metallicities in Figure 5.
We report a mean radial velocity of for Sgr dSph, and a mean radial velocity of for M54. Visually, the radial velocities for the Sgr core and M54 are similar, with only difference in their mean. We run a Kolmogorov-Smirnov (KS) test, implemented with Scikit-learn, which determines whether two distributions are significantly one of a kind from each other and are not from the same parent distribution, on our distributions of heliocentric velocity (Press & Teukolsky, 1988). We find a moderate difference between the two distributions, and with a corresponding p-value of 0.004, this indicates that Sgr and M54’s distributions of heliocentric velocity are highly unlikely to come from the same parent sample.
We report a range of [Fe/H] of to dex for the Sgr core. The Sgr core has a wide-range of [Fe/H] (seen visually in the bottom left CMD of Figure 3 with a high of nearly solar [Fe/H], to a low of dex, matching the range found in Siegel et al. (2007); Alfaro-Cuello et al. (2019). In the middle panel of Figure 5, APOGEE DR17 Sgr members are identified in blue, which form a main peak around dex, then an extended tail which peaks at a lower stellar density at dex. While not on a one-to-one scale, this distribution mimics the metallicity distribution found within Minelli et al. (2023). Hayes et al. (2020) finds a dominant peak of using APOGEE DR16, with a trailing end into the metal-poor regime down to a minimum of , which matches our findings using APOGEE DR17.
M54 has a [Fe/H] range of to dex. We find two distinct peaks within the metallicity distribution, suggesting the existence of two distinct populations with M54, unlike the gradient seen in the Sgr core (see Figure 5.) Boecker et al. (2020), using the Multi-Unit Spectroscopic Explorer (MUSE) spectrum, finds two dominant populations; one is old (8-14 Gyr) with a and a younger (1 Gyr) with a . These measurements from Boecker et al. (2020) appear to line up with our samples’ dominant populations within our M54 members.
Both Sgr and M54 extend to similar ranges within total proper motion , and initially appear to share the same distribution (bottommost panel, Figure 5), albeit that Sgr has more spread than M54. This is confirmed by our measured FWHMs from Section III.3, where Sgr has a , while M54 has a . We do not see any major-minor axis asymmetry with these measurements. The width of these peaks suggest that the internal kinematics within each system are similar but still distinctly different. We run another KS-test to ensure that our distributions of proper motion are not from the same parent sample. We report a minor difference of 4.2% between the two distributions, but with a corresponding p-value of 0.0002, this difference is statistically significant, meaning M54 and Sgr’s proper motion distributions are highly unlikely to come from the same parent sample.
@@
The stellar density contours present in the bottom panel of Figure 1 instead showcase a concentration in the central area of the Sgr core near M54, implying decoupling between the motion of the Sgr core its central concentration peak, possibly indicating that M54 remains mainly un-disrupted despite being embedded in the Sgr core.
With the similar near-solar metallicity populations of both the Sgr and M54 system, it’s possible ( and that this dominant near-solar stellar population within M54 is actually a subpopulation of Sgr if caused by the initial infall of M54, thus should be referred to as such. These stars should then be located near the cluster outskirts of M54. The bottom right panel of Figure 6 clearly shows an even distribution of APOGEE DR17 sources (pink stars), meaning our near-solar stellar population does not just come from the outskirt of the cluster. Sgr on the other hand, most APOGEE DR17 sources are centered nearest to where M54 has been removed from the sample, but there is still some scatter in other parts of the core.
@@
Evidence supporting the theory that M54 formed independently of the Sgr core comes from their respective red giant branch from their CMDs made from our sample in Figure 3. In the bottom panels, where the RGB and AGB branches are indicated with APOGEE DR17 [Fe/H], the Sgr core displays a rough gradient along its RGB and corresponding AGB branches, marked by three different metallicities. This lines up with the three pericentric interactions with the core of the Milky Way that kick-started star formation in both galaxies (de Boer et al., 2015; Ruiz-Lara et al., 2020). The CMD of M54 lacks these three distinct RGB and AGB branches, instead bearing two distinct branches as found in Figure 2 of Mucciarelli et al. (2017). This suggests that M54 only went through one or two pericentric passages, enriching the cluster, kick-starting star formation, implying M54 infell during the first passage of Sgr , which is reflected in the findings of two dominant stellar populations in Boecker et al. (2020). Similarly, the density contours present in the bottom right panel of Figure 1 and the jump from the uniform center to a scattered value outside the first density contour lines with the central density cusp found in Monaco et al. (2005) validate our sample further.
@@
The most recent paper for a Sgr core membership selection was Minelli et al. (2023), where they identified 450 member stars in Early Data Release 3 (EDR3). Cross-matching with our sample, we recover 447 stars from Minelli’s sample. From those 447 stars, none have [Fe/H] measurements from APOGEE DR17, and there are no DR3 [M/H] measurements using General Stellar Parametrizer from Spectroscopy (GSP-Spec) for our sample. There are [M/H] measurements for these 427 within DR3 General Stellar Parametrizer from Photometry (GSP-Phot), however these values exhibit substantial biases compared to literature values and are only useful at a qualitative level especially for red giant branch stars with low quality parallax measurements (Andrae et al., 2023). Other attempts to gain [Fe/H] measurements from The Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) and The GALactic Archaeology with HERMES (GALAH) fail to capture Sgr dSph core and M54 member stars within our sample, unfortunately not allowing us a direct comparison of Minelli et al., 2023 [Fe/H] measurements taken from the Fibre Large Array Multi Element Spectrograph (FLAMES) on the Very Large Telescope (VLT).
VI Conclusion
@@
We demonstrate the strength of DR3 astrometry combined with APOGEE DR17 spectra to separate the Sgr core with M54, along with a multi-Gaussian fit to reduce MW contaminants for systems near the disk. We have a membership catalog of 144596 total stars in the Sgr core and 2638 in M54.
We measure the distance to both the Sgr core and M54 using red clump distancing methods from a cross-matched sample with 2MASS, finding a heliocentric distance of kpc for the Sgr Core and kpc for M54. Our distance measurements agree with prior literature but are more precise due to the larger member sample identified in our work and a new RC distance determination method.
We analyze the motion, density contours, and [Fe/H] of the Sgr core and M54 to understand whether M54 formed independently and was captured during a passage, or is the Sgr nucleus. Preliminary distance measurements and the dynamics of our sample imply they co-exist, but the chemical abundance history suggest the systems formed independently.
@@ VII Acknowledgments
E.K.H.T.T. would like to additionally extend gratitude to Dr. Robyn Sanderson from the University of Pennsylvania, Dr. Sukanya Chakrabarti from the University of Alabama, Huntsville, Dr. Marc Hon from the Massachusetts Institute of Technology, Dr. Andrea Dobson from Whitman College, Dr. Matthew Shetrone from the University of Californa, Santa Cruz, Aldo Sepulveda from the University of Texas at Austin School of Law, and Dr. Keith Hawkins from the University of Texas at Austin and Dr. Eugene Vasiliev from the University of Surrey for their insightful conversations and feedback on this work.
@@
We also thank the reviewer for their helpful feedback and comments to improve this work.
@@
Appendix A Gaia DR3 ADQL Queries
The following DR AQDL query produces our spatial sample for the Sgr dSph core in full.
SELECT
@@ Appendix A Gaia DR3 ADQL Queries
Appendix B Flattened Coordinates
Below we showcase the equations to compute the flattened coordinates for the Sgr core and account of the elliptical shape of the core.The basic principle of the flattened coordinates is to set up a new reference frame centered on the object of study. We also provide definitions for each of the variables: right ascension of the Sgr core (), declination of the Sgr core (), tangent-plane coordinate along the increasing right-ascension direction (), tangent-plane coordinate along the increasing declination direction (), position angle of the major axis of the Sgr core (), rotation angle used to align the coordinate system with the major axis of the Sgr core (), rotated tangent-plane coordinate aligned with the major axis of the Sgr core (), rotated tangent-plane coordinate aligned with the minor axis (), ellipticity (), and elliptical radius from the center of the Sgr core in the flattened coordinate system ().
Appendix C Polygon Coordinates
| Evolutionary Feature | Polygon Coordinates |
| Young | [0.7, 17.8], [-0.5, 17.8], [0.2, 15], [0.7, 15], [0.7, 17.8] |
| [2.6, 15], [2.0, 16.2], [1.6, 17.8], [3.0, 19.5] | |
| AGB | [1.2, 15], [1.2, 13], [1.5, 13], [4.0, 13], [4.0, 15], [1.2, 15] |
| BL | [1, 16], [0.7, 15], [0.5, 12], [1.2, 12], [1.2, 13], [1.2, 15] |
| RC | [1.1, 18.1], [1.1, 17.8], [1.4, 17.8], [1.4, 18.1], [1.1, 18.1] |
| Evolutionary Feature | Polygon Coordinates |
| MS | [-0.5, 20.5], [-0.5, 20], [2.5, 20], [2.5, 20.5], [-0.5, 20.5] |
| [2.6, 15], [2, 16.2], [1.6, 17.8], [3, 19.5] | |
| AGB | [1.2, 15], [1.2, 14], [1.5, 14], [4.4, 14], [4.4, 15], [1.2, 15] |
| RC | [1.1, 18.1], [1.1, 17.8], [1.5, 17.8], [1.5, 18.1], [1.1, 18.1] |
Appendix D Gaia BP-RP Contamination
Appendix E Gaia DR3 Isochrones